OpenStudy (anonymous):

solve sin2x=cosx for exact solutions. 0<=x<=2pi

OpenStudy (anonymous):

0<=x<=2 pi

OpenStudy (jamesj):

Well, what's an expression for sin 2x ?

OpenStudy (anonymous):

2sinxcosx

OpenStudy (jamesj):

Right. So if 2 sin x . cos x = cos x we can reduce that to sin x = 1/2 Now you just need to find x in the given range for which this is true.

myininaya (myininaya):

\[2\cos(x)\sin(x)-\cos(x)=0 =>\cos(x)(2\sin(x)-1)=0\]

myininaya (myininaya):

\[\cos(x) =0 \text{ or} \sin(x)=\frac{1}{2}\]

OpenStudy (jamesj):

true ... was getting there, but he's gone, so I'm glad we've got it on the record.

myininaya (myininaya):

lol yes

OpenStudy (jamesj):

*sigh* hate it when they don't stay and engage with you.

OpenStudy (anonymous):

i'm sorry, my computer shut down on me. but thanks a lot for the help

OpenStudy (jamesj):

ok